OpenStudy (anonymous):

transform each polar equation to an equation in rectanglular coordinates and identify its shape. a. θ=1.34 radians b. r=tanθsecθ

OpenStudy (anonymous):

x=tan(1.34) sec(1.34) cos(1.34) y=tan(1.34) sec(1.34) sin(1.34)

OpenStudy (dumbcow):

in converting polar coordinates to rectangular coordinates, use the following: \[x = r*cos(\theta)\] \[y = r*\sin(\theta)\] \[\theta = \tan^{-1}(\frac{y}{x})\] Now given is \[r = \tan(\theta)*\sec(\theta)\] we also know that \[r = \frac{x}{\cos(\theta)} = \frac{y}{\sin(\theta)}\] By substitution \[\frac{x}{\cos(\theta)} = \tan(\theta)*\sec(\theta)\] \[x = \tan(\theta)\] By substitution \[\frac{y}{\sin(\theta)} = \frac{x}{\cos(\theta)}\] \[y = x*\tan(\theta) =x^{2}\] Therefore the rectangular form of this equation is y=x^2 which is a parabola. when \[\theta = 1.34\] \[r = \tan(1.34)*\sec(1.34) = 18.6\] \[x = \tan(1.34) = 4.256\] \[y = x^{2} = 4.256^{2} = 18.11\] The polar coordinates of the point are (1.34,18.06) The rectangular coordinates are (4.256,18.11)